Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Permutaions-five distinguishable balls

Expert replies
by parveen110 » Sun Apr 27, 2014 5:00 am
Five distinguishable balls are placed into three distinguishable boxes. What is the probablity that atleast one box is empty?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon Apr 28, 2014 3:13 am
parveen110 wrote:Five distinguishable balls are placed into three distinguishable boxes. What is the probablity that atleast one box is empty?
P(at least 1 box is empty) = 1 - P(no box is empty).

Total ways to distribute the balls:
Number of options for the 1st ball = 3. (Any of the 3 boxes.)
Number of options for the 2nd ball = 3. (Any of the 3 boxes.)
Number of options for the 3rd ball = 3. (Any of the 3 boxes.)
Number of options for the 4th ball = 3. (Any of the 3 boxes.)
Number of options for the 5th ball = 3. (Any of the 3 boxes.)
To combine these options, we multiply:
3*3*3*3*3 = 3�.

There are 2 cases in which no box will be empty.

Case 1: One special box gets 3 balls, the other 2 boxes get 1 ball each
Number of box options for the special box = 3. (Any of the 3 boxes.)
From the 5 balls, the number of ways to choose 3 balls for the special box = 5C3 = (5*4*3)/(3*2*1) = 10.
Number of ball options for the next box = 2. (Either of the 2 remaining balls.)
Number of ball options for the last box = 1. (Only 1 ball left.)
To combine these options, we multiply:
3*10*2*1 = 60.

Case 2: One special box gets 1 ball, the other 2 boxes get 2 balls each
Number of box options for the special box = 3. (Any of the 3 boxes.)
Number of ball options for the special box = 5. (Any of the 5 balls.)
From the 4 remaining balls, the number of ways to choose 2 balls for the next box = 4C2 = (4*3)/(2*1) = 6.
From the 2 remaining balls, the number of ways to choose 2 balls for the last box = 2C2 = (2*3)/(2*1) = 1.
To combine these options, we multiply:
3*5*6*1 = 90.

Thus:
P(no box is empty) = (60+90)/3� = 50/81.
Thus:
P(at least 1 box is empty) = 1 - 50/81 = [spoiler]31/81[/spoiler].
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Matt@VeritasPrep » Tue Apr 29, 2014 2:21 pm
Here's an easier but more treacherous way:

Suppose we want the probability that Box 1 is empty. When we place the first ball, the probability is (2/3) that goes in either Box 2 or Box 3. Since we want each of the five balls to go in Box 2 or Box 3, the probability is really (2/3)�.

Continuing like this, the probability that Box 2 is empty is (2/3)� and the probability that Box 3 is empty is (2/3)�.

However, we've overcounted! The probability that Boxes 1 and 2 are both empty appears TWICE (in the probability for Box 1 and the probability for Box 2). So we have to subtract that ONCE. The probability that Boxes 1 and 2 are both empty is (1/3)�, as all the balls would have to go in Box 3.

The probabilities for Box 2 and Box 3 both empty and Box 1 and Box 3 both empty are likewise (1/3)�.

The probability that all three boxes are empty is 0, as we have to pull the balls in SOME box.

Hence we have

Prob(1 empty) + Prob(2 empty) + Prob(3 empty) - Prob(1+2 empty) - Prob(1+3 empty) - Prob(2+3 empty) + Prob(all empty) =

(2/3)� + (2/3)� + (2/3)� - (1/3)� - (1/3)� - (1/3)� + 0 =

3 * ((2/3)� - (1/3)�) = 31/81.
Join the discussion